Pleas Solve
4/(x+3) + 3/(2x-1) =1
4(2x-1)/(x+3)(2x-1) + 3(x+3)/(x+3)(2x-1)

Answer by stanbon(75887)

4/(x+3) + 3/(2x-1) =1
Multiply thru by (x+3)(2x-1) to get:
4(2x-1) + 3(x+3) = (x+3)(2x-1)
8x-4+3x+9 = 2x^2+5x-3
2x^2 -6x -8 = 0
x^2 -3x -4 = 0
Factor:
(x-4)(x+1) = 0
x = 4 or x = -1
=============================

4(2x-1)/(x+3)(2x-1) + 3(x+3)/(x+3)(2x-1)
----
The denominators are the same, so just add the numerators
over the common denominator.
---
= [8x-4+3x+9]/[(x+3)(2x-1)]
---
= [12x+5] / [(x+3)(2x-1)]
=================
Cheers,
Stan H.
===================